OpenStudy (anonymous):

Helpp! Consider the sequence 2,6,18,54... Let n = the term in the sequence let a (n)= the vaule of the nth term of the sequence( see table below) what is the common ratio of the sequence? i think the common ratio is .. times 3 drawling graph

jimthompson5910 (jim_thompson5910):

you said the common ratio was 3 correct?

jimthompson5910 (jim_thompson5910):

if you said that, then you are right because 6/2 = 3 18/6 = 3 54/18 = 3

OpenStudy (anonymous):

yeah the common ratio is 3 but i have a graph that needs filled out and im lost i will post it

jimthompson5910 (jim_thompson5910):

ok

OpenStudy (anonymous):



OpenStudy (anonymous):

What is the realtionship between the exponet ( that was filled in in the graph) and te vaule of n?

OpenStudy (anonymous):

1 more. Complete the stament A(n) = 2x3 this is a fuction to describe the sequence

jimthompson5910 (jim_thompson5910):

notice how when n = 1, A(n) = 2 so if you're multiplying by 3 each time, then the exponent would have to be 0 for the product to be 1 this is because you have (first term)*(common ratio)^exponent

jimthompson5910 (jim_thompson5910):

the first term is 1 the common ratio is 3 so the exponent must be 0 for the product to be 1

jimthompson5910 (jim_thompson5910):

oh sry, the first term is 2, not 1

jimthompson5910 (jim_thompson5910):

so when n = 1, the exponent is 0 This means that the relationship is exponent = n-1 which explains why the formula is a(n) = a*(r)^(n-1)

OpenStudy (anonymous):

im confused :( so all exponest of the graph is 0?

OpenStudy (anonymous):

What is the realtionship between the exponet ( that was filled in in the graph) and te vaule of n? what is this?

jimthompson5910 (jim_thompson5910):

no, that's just for the first term the first term has an exponent of 0

OpenStudy (anonymous):

okay so how do i know the rest?

jimthompson5910 (jim_thompson5910):

for each term after that, the exponent is one less than the term n ex: in the second term, n = 2 but the exponent is 1 in the third term, n = 3 but the exponent is 2 etc etc

jimthompson5910 (jim_thompson5910):

so in general, if the term is say k , then the exponent is k-1 replace the k with n so if you're at the nth term, the exponent is n-1

OpenStudy (anonymous):

What is the realtionship between the exponet ( that was filled in in the graph) and te vaule of n? Complete the stament A(n) = 2x3 this is a fuction to describe the sequence and how do i figure out these?

jimthompson5910 (jim_thompson5910):

so it would be A(n) = 2 x 3^(n-1)

OpenStudy (anonymous):

for the first one?

jimthompson5910 (jim_thompson5910):

What is the realtionship between the exponet ( that was filled in in the graph) and te vaule of n? the term is n, then the exponent is n-1

OpenStudy (anonymous):

okay and then A(n) = 2 x 3^(n-1) this would be the coplete statment?

jimthompson5910 (jim_thompson5910):

yes exactly

OpenStudy (anonymous):

thank you sooo much

jimthompson5910 (jim_thompson5910):

np